# Subset Balancing and Generalized Subset Sum via Lattices

### 06 Apr 2026 v1

@@ 4 Average Case Algorithms for Generalized Subset Sum
In Section˜4.1, we first show that the worst case of Generalized Subset Sum with C=[d:d]C=[-d:d] could reduce to $\text{CVP}_{\infty}$. However, as far as we know, there is no single exponential time algorithm for $\text{CVP}_{\infty}$. Thus we consider the average-case that $\mathbf{x}$ is uniformly random over $0:M-1]^{n}$ and $|\tau|=o(Mn)$ as [6], where we show we can solve it with single exponential time. And we also explore the case where $C=[\pm d]$ in Section˜4.2.
4.1 Generalized Subset Sum with C=[d:d]C=[-d:d]
@@
We first show that the worst case of Generalized Subset Sum with C=[d:d]C=[-d:d] could reduce to $\text{CVP}_{\infty}$.
@@
Given any $d,n>0$, Generalized Subset Sum on C=[d:d]C=[-d:d] can be solved by a single call of a $\text{CVP}_{\infty}$ oracle in dimension $n+1$.
@@
Make a single call to a $\mathrm{CVP}_{\infty}$ oracle on input $(\mathcal{L}_{\alpha,q},\mathbf{t})$, and let $\mathbf{v}=(v_{0},v_{1},\dots,v_{n})\in\mathcal{L}_{\alpha,q}$ be a closest lattice vector returned by the oracle.
@@
so $\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L}_{\alpha,q})\leq d$. Therefore any correct $\mathrm{CVP}_{\infty}$ oracle must return some $\mathbf{v}\in\mathcal{L}_{\alpha,q}$ with $\|\mathbf{v}-\mathbf{t}\|_{\infty}=\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L}_{\alpha,q})\leq d$, and the algorithm will accept. Moreover, by Lemma 4.1, this accepted $\mathbf{v}$ corresponds to coefficients $(v_{1},\dots,v_{n})$ satisfying $\sum_{i}v_{i}x_{i}=\tau$ and $|v_{i}|\leq d$, so the output is valid.
@@
If no such $\mathbf{c}$ exists, then Lemma 4.1 implies $\mathcal{L}_{\alpha,q}\cap(\mathbf{t}+d\,\mathbf{B}_{\infty}^{n+1})=\varnothing$, i.e. $\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L}_{\alpha,q})>d$. Thus the closest vector $\mathbf{v}$ must satisfy $\|\mathbf{v}-\mathbf{t}\|_{\infty}>d$, and the algorithm reports no solution exists.
@@
This uses a single $\mathrm{CVP}_{\infty}$ call in dimension $n+1$. ∎
However, as far as we know, there is no single exponential time algorithm for $\text{CVP}_{\infty}$. Thus we consider the average-case that C=[d:d]C=[-d:d], $\mathbf{x}$ is uniformly random over $0:M-1]^{n}$.
@@
Proof of Theorem˜1.2.
Just directly apply the proof of Theorem˜4.8.∎
@@ 4.2 Generalized Subset Sum with $C=[\pm d]$
In the case when $C=[\pm d]$, it seems much harder for lattice algorithm. In previous seting C=[d:d]C=[-d:d], the solution has a one-to-one correspondence with the shortest (or closest) vector. However, in the situation that $C=[\pm d]$, the shortest vector may corresponding to an invalid solution.
To overcome this, the idea is to enumerate all lattice vectors closed to the target vector using Lemma 2.12. Then we filter out all the vectors that met the conditions.
@@
For $d=1$ and $C=\{-1,1\}$, we use brute-force enumeration, otherwise, we apply Theorem˜4.10.∎

### 24 Apr 2026 v2

@@ 4 Worst Case Algorithms for Generalized Subset Sum
In Section˜4.1, we first show that the worst case of Generalized Subset Sum with C=[d:d]C=[-d:d] could reduce to $\text{CVP}_{\infty}$. However, as far as we know, there is no single exponential time algorithm for $\text{CVP}_{\infty}$.
We observe that the distance from the target to any lattice point can take only integer values, which lets us replace an exact $\mathrm{CVP}_{\infty}$ oracle by an $\left(1+\frac{1}{d}-\varepsilon\right)$ approximate one for the C=[d:d]C=[-d:d] case. Then we apply the oracle proposed by Eisenbrand, Hähnle, and Niemeier [11] in time $2^{O(n)}\cdot(\log(1/\varepsilon))^{O(n)}$, yielding the overall complexity $2^{O(n\log\log d)}$.
4.1 Worst Case Generalized Subset Sum with C=[d:d]C=[-d:d]
@@
It is easy to show that the worst case of Generalized Subset Sum with C=[d:d]C=[-d:d] could reduce to $\text{CVP}_{\infty}$. However, as far as we know, there is no single exponential time algorithm for $\text{CVP}_{\infty}$.
To overcome this for the worst case GSS, our key observation is that the distances are discrete integers. We can exploit this gap: if a solution exists within distance $d$, then we conclude $\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L})\leq d$; otherwise no solution exists, then $\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L})\geq d+1$. This means we only need an approximate $\text{CVP}_{\infty}$ oracle with approximation factor $\left(1+\frac{1}{d}-\varepsilon\right)$ for arbitrarily small $\varepsilon>0$.
@@
Given any $d,n>0$, Generalized Subset Sum on C=[d:d]C=[-d:d] can be solved by a single call of a $(1+1/d-\varepsilon)$-approximate $\text{CVP}_{\infty}$ oracle in dimension $n+1$ for any $0<\varepsilon<1/d$.
@@
Make a single call to a $(1+1/d-\varepsilon)$-approximate $\mathrm{CVP}_{\infty}$ oracle on input $(\mathcal{L}_{\alpha,q},\mathbf{t})$, and let $\mathbf{v}=(v_{0},v_{1},\dots,v_{n})\in\mathcal{L}_{\alpha,q}$ be the lattice vector returned by the oracle.
@@
so $\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L}_{\alpha,q})\leq d$. Therefore the $1+1/d-\varepsilon$-approximate $\mathrm{CVP}_{\infty}$ oracle must return some $\mathbf{v}\in\mathcal{L}_{\alpha,q}$ with
$\|\mathbf{v}-\mathbf{t}\|_{\infty}\leq\left(1+\frac{1}{d}-\varepsilon\right)\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L}_{\alpha,q})<d+1\Longrightarrow\|\mathbf{v}-\mathbf{t}\|_{\infty}\leq d,$
and the algorithm will accept. Moreover, by Lemma 4.1, this accepted $\mathbf{v}$ corresponds to coefficients $(v_{1},\dots,v_{n})$ satisfying $\sum_{i}v_{i}x_{i}=\tau$ and $|v_{i}|\leq d$, so the output is valid.
@@
If no such $\mathbf{c}$ exists, then Lemma 4.1 implies $\mathcal{L}_{\alpha,q}\cap(\mathbf{t}+d\,\mathbf{B}_{\infty}^{n+1})=\varnothing$, i.e. $\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L}_{\alpha,q})\geq d+1$. Thus even the closest vector $\mathbf{v}$ must satisfy $\|\mathbf{v}-\mathbf{t}\|_{\infty}>d$, and the algorithm reports no solution exists.
@@
This uses a $(1+1/d-\varepsilon)$-approximate $\mathrm{CVP}_{\infty}$ oracle in dimension $n+1$. ∎
Eisenbrand, Hähnle, and Niemeier [11] gave a randomized algorithm for $(1+\varepsilon)$-approximate CVP that runs in time
$2^{O(n)}\cdot(\log(1/\varepsilon))^{O(n)},$
where $n$ is the lattice dimension. Their algorithm works by constructing a $(2,\varepsilon)$-covering of the $\ell_{\infty}$ ball using $O((1+\log(1/\varepsilon))^{n})$ parallelepipeds, and invoking a $2$-approximate CVP oracle for each.
Derandomization. The covering construction in [11] is deterministic based on a coordinate-wise binary partition of the cube. Their original algorithm invokes a randomized $2$-approximate CVP oracle by [4] as a subroutine. We instantiate this oracle with the deterministic $2^{O(n)}$-time algorithm of Dadush and Kun [8] by setting $\varepsilon=1$ in their algorithm. This yields a fully deterministic algorithm.
Proposition 4.3 (Deterministic Approximate CVP [11]).
There is a deterministic algorithm for $(1+\varepsilon)$-approximate CVP in $\ell_{\infty}$ norm in time
$2^{O(n)}(\log(1/\varepsilon))^{O(n)}.$
The proof can be found at appendix Appendix˜A.
Theorem 4.4 (Algorithm for Worst-Case GSS with C=[d:d]C=[-d:d]).
Given any $d>2,n\in\mathbb{N}$ and let C=[d:d]C=[-d:d], there is a deterministic algorithm for Generalized Subset Sum in running time
$2^{O(n\log\log d)}.$
Proof.
Combining Theorem 4.2 and Proposition 4.3, for $d>2,\log\log d>0$, set the approximate factor $1+1/(2d)$ completes the proof. ∎
4.2 Worst Case Generalized Subset Sum with $C=[\pm d]$
In the case when $C=[\pm d]$, it seems much harder for lattice algorithm. In previous seting C=[d:d]C=[-d:d], the solution has a one-to-one correspondence with the shortest (or closest) vector. However, in the situation that $C=[\pm d]$, the closest vector may corresponding to an invalid solution.
Warm-up: from C=[d:d]C=[-d:d] to a general interval $C=[a:b]$.
Before proving the reduction formally, we note that essentially nothing changes when we replace the symmetric range [d:d][-d:d] by an arbitrary integer interval $[a:b]$: we keep the same lattice $\mathcal{L}_{\alpha,q}$ and only shift the target to recenter the allowed coefficient window. Concretely, letting $\mu=(a+b)/2$ and r=(ba)/2r=(b-a)/2, we replace $\mathbf{t}=(\alpha\tau,0,\ldots,0)$ by
$\mathbf{t}=(\alpha\tau,\mu,\ldots,\mu),$
so that the constraint $c_{i}\in[a:b]$ is captured by the cube condition $\|\,(c_{1},\ldots,c_{n})-(\mu,\ldots,\mu)\,\|_{\infty}\leq r$. With appropriate choices of $\alpha,q$, this yields
$\mathcal{L}_{\alpha,q}\cap(\mathbf{t}+r~\mathbf{B}_{\infty}^{n+1})=\left\{(\alpha\tau,c_{1},\dots,c_{n})\Bigg|\sum_{i=1}^{n}c_{i}x_{i}=\tau,a\leq c_{i}\leq b\right\},$
and therefore feasibility reduces to a single $\mathrm{CVP}_{\infty}$ instance. Moreover, since the $\ell_{\infty}$-distance is discrete, we get a gap:
$\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L}_{\alpha,q})\leq r\quad\text{(YES-instance)},\qquad\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L}_{\alpha,q})\geq r+1\quad\text{(NO-instance)}.$
Hence a $\bigl(\tfrac{r+1}{r}-\varepsilon\bigr)=\bigl(\tfrac{b-a+2}{b-a}-\varepsilon\bigr)$-approximate $\mathrm{CVP}_{\infty}$ oracle suffices.
From one interval to $C=[\pm d]$ via $2^{n}$ shifted targets.
When $C=[\pm d]=[-d,-1]\cup[1,d]$, we view $C^{n}$ as a union of boxes aligned with the $2^{n}$ axis, one for each choice of coordinate sign. For each sign pattern $\mathbf{s}\in\{\pm 1\}^{n}$, we apply the interval routine to the corresponding box by using a target whose last $n$ coordinates are shifted to the appropriate midpoint:
$\mathbf{t}_{\mathbf{s}}=\bigl(\alpha\tau,\,s_{1}\cdot\tfrac{d+1}{2},\,\ldots,\,s_{n}\cdot\tfrac{d+1}{2}\bigr),\qquad r=\tfrac{d-1}{2},$
while keeping $\mathcal{L}_{\alpha,q}$ unchanged. Thus worst-case GSS for $C=[\pm d]$ is solved by enumerating $\mathbf{s}$ and making $2^{n}$ calls to an approximate $\mathrm{CVP}_{\infty}$ oracle; the same distance-gap argument shows that approximation factor
$\frac{r+1}{r}-\varepsilon\;=\;\frac{d+1}{d-1}-\varepsilon$
is enough for each call. Plugging in the known deterministic $(1+\varepsilon)$-approximate $\mathrm{CVP}_{\infty}$ algorithm (Proposition 4.3) then gives overall time $2^{O(n\log\log d)}$ (with the extra $2^{n}$ factor absorbed into the exponent).
Theorem 4.5 (Algorithm for Worst-Case GSS with $C=[\pm d]$).
Given any $d>3,n\in\mathbb{N}$ and let $C=[\pm d]$, there is a deterministic algorithm for Generalized Subset Sum in running time
$2^{O(n\log\log d)}.$
Proof.
Let r=(d1)/2r=(d-1)/2. Choose the same parameters $\alpha=d+1,q=d\sum_{i=1}^{n}|x_{i}|+|\tau|+1$, so that $\alpha>d$ and $q>d\sum_{i=1}^{n}|x_{i}|+|\tau|$. Construct the lattice $\mathcal{L}_{\alpha,q}$ with basis $\mathbf{B}_{\alpha,q}$ as in Section˜4.1.
Algorithm.
For each sign pattern $\mathbf{s}=(s_{1},\ldots,s_{n})\in\{\pm 1\}^{n}$, define the target
$\mathbf{t}_{\mathbf{s}}=\bigl(\alpha\tau,\;s_{1}\cdot\tfrac{d+1}{2},\;\ldots,\;s_{n}\cdot\tfrac{d+1}{2}\bigr)\in\mathbb{R}^{n+1}.$
Invoke a $\gamma$-approximate $\mathrm{CVP}_{\infty}$ oracle on $(\mathcal{L}_{\alpha,q},\mathbf{t}_{\mathbf{s}})$, where we will set
$\gamma=1+\frac{1}{d-1}=1+\frac{1}{2r}.$
Let $\mathbf{v}=(v_{0},\ldots,v_{n})\in\mathcal{L}_{\alpha,q}$ be the returned lattice vector. If $\|\mathbf{v}-\mathbf{t}_{\mathbf{s}}\|_{\infty}\leq r$, output $(c_{1},\ldots,c_{n}):=(v_{1},\ldots,v_{n})$ and halt. If no $\mathbf{s}$ yields acceptance, reports that no solution exists.
A sign-pattern version of Lemma 4.1.
Fix $\mathbf{s}\in\{\pm 1\}^{n}$. The proof of Lemma 4.1 extends verbatim after shifting the last $n$ target coordinates from 0 to $s_{i}\cdot\tfrac{d+1}{2}$: one shows that
$\mathcal{L}_{\alpha,q}\cap\bigl(\mathbf{t}_{\mathbf{s}}+r\,\mathbf{B}_{\infty}^{n+1}\bigr)=\Bigl\{(\alpha\tau,c_{1},\ldots,c_{n})\;\Bigm|\;\sum_{i=1}^{n}c_{i}x_{i}=\tau,\ \ s_{i}c_{i}\in[1:d]\Bigr\}.$
In particular, $\mathcal{L}_{\alpha,q}\cap(\mathbf{t}_{\mathbf{s}}+r\,\mathbf{B}_{\infty}^{n+1})\neq\varnothing$ if and only if there exists a feasible GSS solution whose coordinate-wise signs are given by $\mathbf{s}$.
Distance gap.
Fix $\mathbf{s}\in\{\pm 1\}^{n}$ and write $\mathbf{t}_{\mathbf{s}}=(\alpha\tau,\,s_{1}\frac{d+1}{2},\ldots,s_{n}\frac{d+1}{2})$. For any $\mathbf{v}\in\mathcal{L}_{\alpha,q}$, the first-coordinate difference satisfies
$v_{0}-\alpha\tau\in\alpha\mathbb{Z},$
since $v_{0}=\alpha(qz_{0}+\sum_{i}z_{i}x_{i})$. With our choice $\alpha=d+1$ and r=(d1)/2r=(d-1)/2, we have $r+1=(d+1)/2<\alpha$. Hence
$\|\mathbf{v}-\mathbf{t}_{\mathbf{s}}\|_{\infty}<r+1\ \Longrightarrow\ |v_{0}-\alpha\tau|<\alpha\ \Longrightarrow\ v_{0}=\alpha\tau,$
so the $\ell_{\infty}$-distance is determined entirely by the last $n$ coordinates:
$\|\mathbf{v}-\mathbf{t}_{\mathbf{s}}\|_{\infty}=\max_{i\in[n]}\Bigl|v_{i}-s_{i}\frac{d+1}{2}\Bigr|.$
Now $v_{i}\in\mathbb{Z}$, while $s_{i}\frac{d+1}{2}$ is an integer if $d$ is odd and a half-integer if $d$ is even. Therefore the above maximum ranges over $\mathbb{Z}$ (odd $d$) or $\mathbb{Z}+\tfrac{1}{2}$ (even $d$). Since r=(d1)/2r=(d-1)/2 has the same type, there is no attainable value in $(r,r+1)$, and thus
$\|\mathbf{v}-\mathbf{t}_{\mathbf{s}}\|_{\infty}<r+1\quad\Longrightarrow\quad\|\mathbf{v}-\mathbf{t}_{\mathbf{s}}\|_{\infty}\leq r.$
Correctness and Soundness.
Suppose there exists $\mathbf{c}\in[\pm d]^{n}$ with $\sum_{i}c_{i}x_{i}=\tau$. Let $s_{i}:=\mathrm{sign}(c_{i})\in\{\pm 1\}$. Then $s_{i}c_{i}\in[1:d]$ for all $i$, and by the sign-pattern lemma above, $(\alpha\tau,c_{1},\ldots,c_{n})\in\mathcal{L}_{\alpha,q}\cap(\mathbf{t}_{\mathbf{s}}+r\,\mathbf{B}_{\infty}^{n+1})$, so $\mathrm{dist}_{\infty}(\mathbf{t}_{\mathbf{s}},\mathcal{L}_{\alpha,q})\leq r$. By $\gamma$-approximate $\mathrm{CVP}_{\infty}$,
$\|\mathbf{v}-\mathbf{t}_{\mathbf{s}}\|_{\infty}\leq\gamma\cdot\mathrm{dist}_{\infty}(\mathbf{t}_{\mathbf{s}},\mathcal{L}_{\alpha,q})\leq\Bigl(1+\frac{1}{2r}\Bigr)r=r+\frac{1}{2}<r+1.$
By the gap observation, this implies $\|\mathbf{v}-\mathbf{t}_{\mathbf{s}}\|_{\infty}\leq r$, so the algorithm accepts for this $\mathbf{s}$. Finally, acceptance implies $\mathbf{v}\in\mathcal{L}_{\alpha,q}\cap(\mathbf{t}_{\mathbf{s}}+r\,\mathbf{B}_{\infty}^{n+1})$, and by the sign-pattern lemma the output coefficients $(v_{1},\ldots,v_{n})$ satisfy $\sum_{i}v_{i}x_{i}=\tau$ and $v_{i}\in[\pm d]$.
If the instance has no solution in $\pm d]^{n}$, then for every $\mathbf{s}$ the intersection $\mathcal{L}_{\alpha,q}\cap(\mathbf{t}_{\mathbf{s}}+r\,\mathbf{B}_{\infty}^{n+1})$ is empty. Equivalently, $\mathrm{dist}_{\infty}(\mathbf{t}_{\mathbf{s}},\mathcal{L}_{\alpha,q})>r$, so no lattice vector can satisfy $\|\mathbf{v}-\mathbf{t}_{\mathbf{s}}\|_{\infty}\leq r$. Hence the algorithm never accepts and correctly reports that no solution exists.
Running time.
We make $2^{n}$ approximate $\mathrm{CVP}_{\infty}$ oracle calls in dimension $n+1$. For $d>3$, instantiating the oracle by Proposition 4.3 with $\varepsilon=1/(d-1)$ gives per-call time
$2^{O(n)}\bigl(\log(1/\varepsilon)\bigr)^{O(n)}=2^{O(n)}\bigl(\log(d-1)\bigr)^{O(n)}=2^{O(n\log\log d)}.$
Multiplying by the outer $2^{n}$ factor yields overall time $2^{O(n\log\log d)}$, as claimed. ∎
Proof of Theorem˜1.2.
Combining Theorem˜4.4 and Theorem˜4.5 completes the proof.∎
5 Average Case Algorithms for Generalized Subset Sum
In Section˜4.1, we already show that the worst case of Generalized Subset Sum with C=[d:d]C=[-d:d] could reduce to an approximate $\text{CVP}_{\infty}$ oracle in time $2^{O(n\log\log d)}$. In this section, we consider the average-case that $\mathbf{x}$ is uniformly random over $0:M-1]^{n}$ and $|\tau|=o(Mn)$ as [7], where we show we can solve it with single exponential time without $d$ in Section˜5.1. And we also explore the case where $C=[\pm d]$ in Section˜5.2.
5.1 Average Case Generalized Subset Sum with C=[d:d]C=[-d:d]
In this section we consider the average-case that C=[d:d]C=[-d:d], $\mathbf{x}$ is uniformly random over $0:M-1]^{n}$ as settings in [7].
@@ 5.2 Generalized Subset Sum with $C=[\pm d]$
We resolve the worst-case problem in Section˜4.2 by reducing to approximate $\mathrm{CVP}_{\infty}$ algorithm. In the average setting, we adopt a different approach.
We could also prove a bounded distance promise holds in high probability, then we enumerate all lattice vectors closed to the target vector using Lemma 2.13. Finally we filter out all the vectors that met the conditions.
@@
For the case when C=[d:d]C=[-d:d], applying Theorem˜5.6. For the case when $C=[\pm d]$, we use brute-force enumeration for $d=1$ and $C=\{-1,1\}$, otherwise, we apply Theorem˜5.8.∎
Appendix A Proof of Proposition˜4.3
Proof.
We combine the deterministic $2$-approximation for $\mathrm{CVP}_{\infty}$ obtained from Dadush and Kun [8] with the covering and boosting framework of Eisenbrand, Hähnle, and Niemeier [11].
Step 1: a deterministic $2$-approximation oracle. Dadush and Kun (Theorem 1.1 [8]) give a deterministic algorithm that, for any norm and any $0<\delta\leq 1$, computes a $(1+\delta)$-approximate closest vector in time
$2^{O(n)}\Bigl(1+\frac{1}{\delta}\Bigr)^{\!n}.$
The $\ell_{\infty}$ norm is symmetric, so the algorithm applies directly. Setting $\delta=1$ yields a deterministic $2$-approximation algorithm for $\mathrm{CVP}_{\infty}$, which we denote by $\mathcal{A}_{2}$; its running time is $2^{O(n)}$.
Step 2: $2$-gap oracle from $\mathcal{A}_{2}$. Given a lattice $\mathcal{L}$, a target $\mathbf{t}$ and a distance $D>0$, we scale the problem by $1/D$ so that $D=1$. One call to $\mathcal{A}_{2}$ on $(\mathcal{L},\mathbf{t})$ returns a vector $\mathbf{v}$ with $\|\mathbf{v}-\mathbf{t}\|_{\infty}\leq 2\cdot\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L})$.
  • If $\|\mathbf{v}-\mathbf{t}\|_{\infty}\leq 2$, we have found a lattice point in $\mathbf{t}+[-1,1]^{n}$ (after undoing the scaling).
  • If $\|\mathbf{v}-\mathbf{t}\|_{\infty}>2$, then necessarily $\mathrm{dist}_{\infty}(\mathbf{t},\mathcal{L})>1$, which certifies that $\mathbf{t}+[-1,1]^{n}$ contains no lattice point.
Thus a single call to $\mathcal{A}_{2}$ implements a deterministic $2$-gap oracle as required in [11].
Step 3: boosting to $(1+\varepsilon)$-approximation. Eisenbrand et al. [11] show how to boost any $2$-gap oracle to a $(1+\varepsilon)$-approximation using two completely deterministic subroutines.
  1. 1.
    From $2$-gap to $(1+\varepsilon)$-gap. Theorem 3.2 of [11] reduces the $(1+\varepsilon)$-gap decision problem to at most
    $2^{n}\Bigl(2+\log\frac{1}{\varepsilon}\Bigr)^{n}=2^{O(n)}\Bigl(\log\frac{1}{\varepsilon}\Bigr)^{O(n)}$
    calls of the $2$-gap oracle. The reduction is based on covering the cube $1+\delta,1-\delta]^{n}$ (with $\delta=\varepsilon/(1+\varepsilon)$) by axis-parallel parallelepipeds that, when scaled by $2$, stay inside $1,1]^{n}$ (Theorem 2.6). The covering is built by a coordinate-wise binary partition of the cube and is therefore deterministic. Each parallelepiped induces one instance of the $2$-gap problem, and the encoding length of each instance stays polynomial in $n$, $\log(1/\varepsilon)$ and the original input size $b$.
  2. 2.
    From $(1+\varepsilon)$-gap to $(1+\varepsilon)$-approximation. Theorem 3.4 of [11] gives a deterministic binary search that, given a $(1+\delta)$-gap oracle with $\delta=\min\{\varepsilon/5,1/2\}$, outputs a $(1+\varepsilon)$-approximate closest vector. The search makes
    $O\bigl(\log b+\log n+\log(1/\varepsilon)\bigr)$
    calls to the $(1+\delta)$-gap oracle.
Total complexity. We compose the reductions deterministically:
  • The $2$-gap oracle is implemented by $\mathcal{A}_{2}$.
  • Solving the $(1+\varepsilon)$-gap problem uses $\ell_{1}=2^{O(n)}(\log\frac{1}{\varepsilon})^{O(n)}$ calls to $\mathcal{A}_{2}$.
  • The final binary search uses $\ell_{2}=O(\log b+\log n+\log\frac{1}{\varepsilon})$ calls to the $(1+\varepsilon)$-gap solver.
Hence the overall algorithm calls $\mathcal{A}_{2}$ at most $\ell_{1}\cdot\ell_{2}=2^{O(n)}(\log\frac{1}{\varepsilon})^{O(n)}$ times. Each call takes $2^{O(n)}\mathrm{poly}(b)$ time and the size of every constructed instance remains $\mathrm{poly}(n,b,\log\frac{1}{\varepsilon})$, so the total running time is
[…]
All components are deterministic, thus the complete procedure is a deterministic $(1+\varepsilon)$-approximation for $\mathrm{CVP}_{\infty}$. ∎